# Rethinking Programmed I/O for Fast Devices, Cheap Cores, and Coherent Interconnects

### v1

@@ Abstract.
In this paper we question this wisdom in the light of modern hardware and workloads, particularly in cloud servers. We argue that the assumptions that led to this model are obsolete, and in many use-cases use of programmed I/O, where the CPU explicitly transfers data and control information to and from a device via loads and stores, actually results in a more efficient system.
We quantitatively demonstrate these advantages using three use-cases: fine-grained RPC-style invocation of functions on an accelerator, offloading of operators in a streaming dataflow engine, and a network interface targeting for serverless functions. Moreover, we show that while these advantages are significant over a modern PCI Express (PCIe) peripheral bus, a truly cache-coherent interconnect offers significant additional efficiency gains.
@@ 1. Introduction
There is a current consensus on how CPUs should communicate with modern, high-performance devices like network interface adaptors (NICs), storage controllers, Graphics Processing Units (GPUs), and other computational accelerators: for the most part, this is achieved using DMA and rings of descriptors to decouple CPU and device as much as possible. On modern PCIe interconnects, this consensus results in impressive bandwidth for applications which transfer data in large chunks.
However, in this paper we question this consensus in the light of several factors: (1) important workloads involving irregular access and fine-grained interaction between CPU and device, (2) increasing core counts in servers, (3) data center applications where many machines are executing the same application on most of their cores, and (4) the emergence of cache-coherent peripheral interconnects, particularly those like CXL.cache 3.0 (CXL Consortium, 2024a) in which devices are first-class participants in the coherence protocol.
We argue that these factors make Programmed I/O (PIO), where CPU software performs I/O by directly reading and writing device registers, a better mechanism for data transfer in some important scenarios than one based on DMA.
We use a real hardware platform which implements a coherent interconnect between a server-class CPU and a large FPGA to investigate the trade-offs between DMA using descriptor rings, PIO directly over a PCIe interconnect, and PIO over a cache-coherence protocol.
@@
In the following section, we provide background and motivation for this work, including the factors that led to the descriptor-based DMA model of I/O, and why it is time to question the underlying assumptions in that model. In Section 3 we describe and motivate the experimental hardware platform we use for our evaluation, and also provide calibrating performance comparisons with PC data center servers.
In Section 4, we discuss why true cache-coherent device interconnects are fundamentally different from today’s PCIe platforms, and present a set of efficient protocols for passing messages with low latency between software running on an unmodified CPU, and a smart device which has low-level access to the coherence protocol. In Section 4.5, we go on to present the implementation of some of these protocols on our experimental platform.
@@ 2. Background and Motivation
2.1. Background: hardware/software interaction
In early machines, the CPU interacted with peripheral devices exclusively using programmed I/O (PIO): the device exposed hardware registers mapped into the processor’s address space. Both data transfer and control of the device involved reads and writes by the CPU to these registers.
@@ 2.1. Background: hardware/software interaction
Almost every high-speed device today uses essentially this technique, with minor variations such as how the descriptor queues are formatted, where they are stored, and whether the head and tail of the queue is identified by registers or flags in the descriptors themselves.
This model is optimized for throughput – it is most efficient when a large volume of data must be transferred.
2.2. Trends in interconnects and devices
@@
Recent trends in accelerator design, however, have led to proposed standards for cache coherent peripheral interconnects, and the related idea of extending processor cache coherence to heterogeneous devices: accelerators, networking interfaces, and storage controllers (Stuecheli et al., 2018; CCIX Consortium and others, 2024; CXL Consortium, 2024b; Nvidia, 2024; AMBA® CHI Architecture Specification, 2024). The more far-reaching of these proposals are fully symmetric, i.e. devices are first-class participants in a distributed directory-based cache coherence protocol.
Communication in software between peer modes in a cache-coherent system is typically very different from using descriptor-based DMA. FastForward (Giacomoni et al., 2008a), Barrelfish’ UMP (Baumann et al., 2009a), Concord (Iyer et al., 2023), and Shinjuku (Kaffes et al., 2019) adopt a much more direct approach to sending small messages with low latency, e.g. by exploiting the cache coherence protocol to transfer lines between caches on demand, and local polling to provide synchronization. Recently, cache coherence between NUMA CPUs has been used to simulate communicating between software and a hypothetical cache-coherent NIC (Schuh et al., 2024).
2.3. “New” workloads
Without denying the importance of throughput-oriented data transfer between devices and cores for workloads like current machine learning training and inference, it is worth stepping back and questioning many of the assumptions that led us here, and what other workloads might be a poor fit today for this model.
@@
Any workload involving fine-grained, frequent interaction between the CPU and another device incurs a significant latency penalty when using DMA for data transfer. Many irregular workloads have this property (Gioiosa et al., 2017; Bean, 2016; Kashimata et al., 2019; Zeng and Prasanna, 2020).
@@
Moreover, data center messaging (essentially, network Remote Procedure Calls (RPCs)) exchange many small messages (Karandikar et al., 2021; Lazarev et al., 2021). While end-to-end latency is mostly network propagation time in this case, latency incurred in the end system is a good proxy for resources (such as CPU cycles) wasted before invoking the application code, and in marshaling and sending any reply.
@@
There are also several assumptions underlying the DMA-based communication model which are worth revisiting. In particular, CPU execution and I/O were decoupled because CPU cycles were precious (there being few cores), and the CPU had plenty of different tasks to work on at any time.
@@
This is no longer true. Server CPUs have had many cores for some time, and moreover when running data center applications they are often dedicated to a single application, sometimes with miscellaneous functions moved to a small subset of the cores (Marty et al., 2019).
@@
Another example: for throughput-oriented workloads DMA has evolved to efficiently transfer data to and from main memory without polluting the CPU cache. However, for small message applications, it is important that almost all the data gets into the right CPU cache as quickly as possible.
@@
This leads us to rethink the use of the DMA model, and whether directly involving the CPU in data transfers to and from devices might be a better approach for many scenarios on modern hardware. Others have also suggested this in some cases (Larsen and Ben Lee, 2015; Chaudhari et al., 2017), but we also push the question much further: given a future cache-coherent interconnect, what new ways of using Programmed I/O might become compelling for data center applications.
@@ 3. Experimental platform
We investigate the modern trade-offs between PIO, DMA, and cache-coherent interconnects using the Enzian research computer (Cock et al., 2022). Enzian is useful as a placeholder for future, interoperable, cache-coherent interconnects like CXL.
We use a real hardware platform rather than a simulator. Simulation would be appropriate for detailed quantitative comparisons of known techniques, but in this paper we explore unconventional uses of cache coherence and so want to establish not only that our techniques are performant, but that they can practically be implemented in a real system.
An Enzian machine can be thought of as a two-socket NUMA server, where one socket houses a Marvel Cavium ThunderX-1 CPU, with a Xilinx XCVU9P FPGA in the other.
The CPU is a 48-core ARMv8 processor running at 2.0GHz, and 128GiB of 2133MT/s DDR4 memory spread across 4 memory controllers. Each core has a 32KiB, 32-way associative, write-through, physically indexed, physically tagged, L1 data cache. These are connected to a 16MiB, write-back, 16-way associative, shared L2 last-level cache. Hardware keeps the L1 caches coherent with the L2, and so in ARM architecture parlance the point of coherence is the L1 cache, whereas the point of unification is the L2 cache write-through. The cache-line size is 128 bytes – note that this is double the conventional 64-byte line size.
The CPU and FPGA are connected by the CPU’s native inter-socket cache coherent protocol, the Enzian Coherence Interface (ECI), which Enzian implements on the FPGA. It is a MOESI-like directory based protocol, and the physical address space is statically partitioned between the two NUMA nodes (CPU and FPGA). It uses 24 bidirectional lanes (organized into 2 12-lane links), each running at about 10Gb/s, for a theoretical inter-socket bandwidth of about 30GiB/s.
@@ 3.1. Characterizing the platform
Enzian provides a real hardware platform with server-class performance to evaluate different forms of device communication. However, direct comparisons with modern server platforms is not straightforward: the latter do not offer a full cache coherence protocol between the CPU and devices, but do provide more recent and therefore higher-performance cores, memory systems, and PCIe interconnects.
@@
Therefore, in order to calibrate against a modern PC for both PIO and DMA over PCIe, we present two benchmark comparisons of Enzian with a modern PC (Intel Core i7-7700 3.6GHz Kaby Lake, PCIe Gen.3 x16) with cache line size of 64 bits connected to an AMD Xilinx Virtex UltraScale+ VCU118 card, using the same “write then read” experiment detailed in Section 5.1. The VCU118 uses the same FPGA as Enzian, albeit a slightly slower speed gauge.
3.1.1. DMA performance over PCIe
@@
The first experiment uses DMA, using the Xilinx XDMA IP and associated descriptor-based protocol to transfer data between the CPU and the FPGA over PCIe. Figure 1 shows results for various data transfer sizes, running the Xilinx driver in both interrupt-driven and polled modes.
Figure 1. Comparison of XDMA latency for RPC workload between Enzian and PC
@@
We observe that, using PCIe, exchanging messages using XDMA is about 3 times as fast on the PC server than on Enzian, while the difference between interrupt-driven and polling performance is much less significant. The latency of a single transaction is pretty constant on both platforms up to the PCIe transaction size limit of 4KiB, and then increases.
@@
The difference in performance between the two architectures is due to several factors. For very small transfers, the CPU overhead of setting up descriptors dominates, and an x86 core is simply faster than the slower (and mostly in-order) ThunderX-1 core. In addition, the memory system is somewhat faster on the PC. The factor of 2 difference in PCIe bandwidth between the two platforms does not appear to be a factor in these experiments.
3.1.2. PIO performance over PCIe
Our second calibration experiment involves the CPU itself doing PIO reads and writes over PCIe to the FPGA.
@@ 3.1.2. PIO performance over PCIe
We experimented with several different optimizations in this experiment, in particular the use of write combining and vector instructions. We present the best figures we could obtain on both platforms.
Figure 2. Comparison of PIO latency for RPC workload between Enzian and PC
@@
Figure 2 shows the results. For transaction sizes above 32 bytes, the PC is about twice as fast as Enzian over the PCIe cable between CPU and FPGA. Here, the factor of 2 difference in PCIe bandwidth is responsible for the latency difference, in particular for reads (which are limited to 128 bits on both platforms). Writes, in contrast, are combined and highly pipelined at the PCIe interface.
4. Programmed I/O over coherent interconnects
@@
In this section we explore what programmed I/O might look over coherent interconnects. It is of course possible for a CPU to simply copy data over a coherent interconnect as in conventional programmed I/O. This requires careful management of the processor’s cache hierarchy with explicit cache flush and invalidate instructions, but can be appropriate when the underlying hardware standard itself mandates a cache on the device, as with OpenCAPI 4.0 (Stuecheli et al., 2018; Allison et al., 2020).
@@
That said, here we are more interested in how a coherent interconnect can best be used for low-latency, efficient movement of data between CPU registers and a device, including cases where the conventional notion of the cache as a view over memory contents is not appropriate (such as data arriving over a network).
@@
We scope the question further as follows: we consider the case where the CPU cores, or their Last-Level Cache (LLC), communicate with devices via a MESI-like (Nagarajan et al., 2020), message-based coherence protocol with stable states corresponding to Modified (dirty), Shared (clean), Exclusive (clean), and Invalid, using a distributed cache directory. Note that this means that each address or cache line has a home node, which maintains the directory entry for the line.
@@
Moreover, we assume a symmetric protocol: each device is a first-class participant in the protocol with a subset of physical addresses homed at it, and the device therefore maintains directory state for these lines.
@@
While these assumptions are rare in hardware today, they are true both of the proposed CXL.cache standard (CXL Consortium, 2024a) and the Enzian platform we use for experiments.
4.1. Baseline: naively communicating with coherence
@@
Our starting point is existing techniques which use cache coherence to perform inter-core communication, for example  (Giacomoni et al., 2008a) or Barrelfish UMP  (Baumann et al., 2009a). The basic idea is shown in Figure 3; to software, the sender writes data to a cache line on which the receiver is spinning (reading in a tight loop). Initially, the line is in (say) Shared state in both caches. When the sender writes the line (typically in one go due to the core’s write-combining buffer), it transitions briefly to Modified state, before being fetched by the receiver over the interconnect and reverting to Shared.
@@
This can be highly efficient in software, particularly when deployed as a ring buffer of cache lines, and results in very little interconnect traffic (since spinning is usually on a local copy of the line). One can easily replace one core with a device in this scheme (with suitable state machine) to obtain an I/O protocol. The downside is, of course, the CPU cost of spinning, both in wasted cycles and the inability to switch to another task.
The design of protocols like FastForward is based on the idea that each endpoint can only manipulate its local cache using software instructions: loads, stores, etc. Moreover, it is assumed (reasonably) that flush or invalidate instructions should be avoided because they might be slow, and might require kernel privilege.
This is not the case when one of the endpoints is a device with message-level access to the cache coherence protocol, and possibly maintaining its own directory state.
4.2. The implications of low-level access
Unlike software running on a core behind a cache, a device participating in a coherence protocol has visibility over more state and events, and can also initiate more operations. Note that there is no need for a cache per se on the device at all, indeed this is usually inappropriate.
First, the device receives messages from the CPU cache that request lines in Shared or Exclusive state, or which request that the caching state on the be downgraded (e.g. from Exclusive to Shared, or from Shared to Invalid). It is therefore able to react internally to such events as it chooses, as long as it does not violate the protocol.
@@ 4.2. The implications of low-level access
Fourth, unlike a conventional cache, the device does not have to respond immediately to every request from the CPU’s cache. Instead, it can choose to delay the response (blocking the requesting core) until some other operations complete. Care must be taken to respond before any hardware-imposed timeouts, but these are typically generous (hundreds of milliseconds on the ThunderX-1, for example) and can be worked around in software (e.g. by sending a “try again” response within the timeout).
@@
4.3. Communication protocol variants
We present 4 different but similar communication protocols (shown in Figure 4) which vary based on (a) the CPU or the device initiates the message transfer, and (b) the cache lines used are homed on the CPU or device. All provide a synchronous RPC: a transfer of a single cache-line-sized message, followed by a single cache-line response. All use 2 cache lines, which we designate A and B. These protocols can be generalized to larger payloads, and latency will grow linearly with the payload size. Interestingly, they can be made slightly more efficient by using more cache lines, but quickly become complex. We focus on the simple 2-line case in this paper.
The protocols use the same techniques and insights, and are best described using the example of a CPU-initiated RPC using addresses homed on the device (4(a)). An implementation of this protocol is used for the evaluation in Section 5.1.
@@ 4.3. Communication protocol variants
    The protocol starts with line B in Exclusive state in the CPU cache, and line A Invalid. To initiate an operation, the CPU writes the request data from its registers into line B, which transitions to Modified. This happens without communication with the device, since B was previously in Exclusive.
@@
    The CPU cache receives line A’s new data and the CPU unblocks, immediately loading the first word of the RPC response into a register.
As described, the resulting state is that A is now Invalid and B is Shared in the CPU cache. However, in some deployed MESI-like protocols (including the Enzian interconnect we use), the device is free to return the data stipulating that it is not Shared but now Exclusive to the CPU. This final feature leaves the cache in the state it started in, with the roles of A and B reversed, ready for the next transaction.
The whole exchange takes 2 round trips on the coherent interconnect.
The other 3 protocols work similarly. For example, 4(b) shows the analogous case where the device initiates the request to the CPU, but still using lines homed on the device. This is used for the NIC experiment in Section 5.3.
@@
4.4. Discussion
A number of concerns might be raised about these protocols in practice; we address them in this section.
First is the issue of timeouts: we are blocking a core’s request for a cache line until a computation has completed. If this triggers a timeout in the CPU’s cache, the result is likely to be a memory fault exception, a machine check, or the processor simply locking up (on the Enzian, it is a machine check). We can work around these issues in practice by a small state machine on the device return a “not ready yet” response in the event of an approaching timeout, causing the software on the issue another request for the other cache line, extending the response time effectively indefinitely without the need for spinning. In practice, timeouts are quite long and the computations we target are short.
@@ 4.4. Discussion
The second issue is more subtle, and concerns possible deadlock due to the device stalling a request until it can issue, and gets a reply back from, another to a different cache line. Such standards as exist for interoperable cache coherent interconnects are silent on what happens in this situation, possibly because it has not occurred until now. The design assumes that, in the platform, transactions on different cache lines can progress independently.

### v2

@@ Abstract.
In this paper we question this wisdom in the light of modern hardware and workloads, particularly in cloud servers. Like others before us, we argue that the assumptions that led to this model are obsolete, and in many use-cases use of Programmed I/O (PIO), where the CPU explicitly transfers data and control information to and from a device via loads and stores, actually results in a more efficient system.
However, unlike others to date, we push this idea further and show, in a real implementation, the gains in average and tail latency for fine-grained communication achievable using an open cache-coherence protocol which exposes cache transitions to a smart device. We show this using three use-cases: fine-grained RPC-style invocation of functions on an accelerator, offloading of operators in a streaming dataflow engine, and a network interface targeting for serverless functions, comparing our use of coherence with both traditional DMA-style interaction and a highly-optimized implementation using PIO over PCI Express (PCIe).
@@ 1. Introduction
Modern interfaces between CPUs and high-performance devices like network interface adaptors (NICs), storage controllers, Graphics Processing Units (GPUs), and other computational accelerators are designed to optimize throughput. Built over a PCIe interconnect, descriptor rings in main memory hold queues of operations and completions, and the device performs DMA transactions to read and write both descriptors and payload data. This design, along with the underlying PCIe interconnect, trades off latency for small transactions in favor of high throughput for large ones.
In this paper we question this consensus in the light of workloads which depend for performance not on bulk throughput, but the cumulative latency of small transactions between CPU and device: closely-coupled hardware accelerators for irregular workloads, or data center Remote Procedure Call (RPC) using small messages. We do this in the context of emerging cache-coherent peripheral interconnects like CXL.cache 3.0 (CXL Consortium, 2024).
The continuing increase in PCIe bandwidth means that, for a given transfer size, the significance of PCIe latency (time to first byte) has become more of a concern, leading to many proposals for reducing the overhead of descriptor management (which we survey in Section 7.
However, we propose a more radical approach, aimed at use-cases where reduced latency is more important than higher maximum throughput: using PIO to Memory-Mapped I/O (MMIO) devices directly, and avoiding DMA, descriptor queues, operation buffering, and interrupts entirely.
We use a real hardware platform which implements a coherent interconnect between a server-class CPU and a large FPGA to investigate the trade-offs between DMA using descriptor rings, PIO directly over a PCIe interconnect, and PIO over a full cache-coherence protocol.
@@
Crucially, not only is median latency considerably lower, but the tail of the latency curve is reduced significantly when using PIO over PCIe, and completely eliminated using the cache coherence protocol.
@@
In the following section, we provide background and motivation for this work, including the factors that led to the descriptor-based DMA model of I/O, and why it is time to question the underlying assumptions in that model. In Section 3 we describe and motivate the experimental hardware platform we use for our evaluation, and also provide calibrating performance comparisons with PC data center servers. We show how to use PIO with PCIe in Section 4.
In Section 5, we discuss why true cache-coherent device interconnects are fundamentally different from today’s PCIe platforms, and present a set of efficient protocols for passing messages with low latency between software running on an unmodified CPU, and a smart device which has low-level access to the coherence protocol. In Section 5, we go on to present the implementation of some of these protocols on our experimental platform.
@@ 2. Background and Motivation
Interconnects and devices:
In early machines, the CPU interacted with peripheral devices exclusively using programmed I/O (PIO): the device exposed hardware registers mapped into the processor’s address space. Both data transfer and control of the device involved reads and writes by the CPU to these registers. This model is also known as memory-mapped I/O (MMIO).
@@
Almost every high-speed device today uses essentially this technique, with minor variations such as how the descriptor queues are formatted, where they are stored, and whether the head and tail of the queue is identified by registers or flags in the descriptors themselves. This model is optimized for throughput – it is most efficient when a large volume of data must be transferred. On modern PCIe interconnects, this consensus results in impressive bandwidth for applications which transfer data in large chunks, up to 64GiB/s for PCIe 5.0 x16.
@@
For the large batch transfers used in GPU-derived programming models (e.g., machine learning workloads), this latency penalty is insignificant when amortized over a large transfer. Indeed, such software has arguably adapted to this model imposed by the hardware (Hooker, 2021).
@@
Recent trends in accelerator design, however, have led to proposed standards for cache coherent peripheral interconnects, and the related idea of extending processor cache coherence to heterogeneous devices: accelerators, networking interfaces, and storage controllers (Stuecheli et al., 2018; CCIX Consortium and others, 2024; CXL Consortium, 2024; NVIDIA, 2024; AMBA® CHI Architecture Specification, 2024). The more far-reaching of these proposals are fully symmetric, i.e. devices are first-class participants in a distributed directory-based cache coherence protocol.
Fine-grained workloads:
@@
While throughput-oriented workloads involving data transfer between devices and cores (e.g., AI training and inference) are hugely important today, we step back in this paper and consider other workloads that might be a poor fit today for this model, because they involve fine-grained, frequent interaction between the CPU and another device. Many irregular workloads have this property (Gioiosa et al., 2017; Bean, 2016; Kashimata et al., 2019; Zeng and Prasanna, 2020).
@@
Moreover, data center messaging (essentially, RPCs) exchange many small messages (Karandikar et al., 2021; Lazarev et al., 2021). While end-to-end latency is mostly network propagation time in this case, latency incurred in the end system is a good proxy for key resources (such as CPU cycles) wasted before invoking the application code, and in marshaling and sending any reply.
All these workloads incur a significant latency penalty when using DMA for data transfer. Moreover, the assumptions which led to the common DMA descriptor-based model do not hold for these workloads.
@@
For example, CPU execution and I/O were decoupled using descriptor rings because CPU cycles were precious (there being few cores), and the CPU had plenty of different tasks to work on at any time.
@@
This is no longer true. Server CPUs now have 10s or 100s of cores, and moreover when running data center applications they are often dedicated to a single application, sometimes with miscellaneous functions moved to a small subset of the cores (Marty et al., 2019). Most of the CPU cores have nothing to do except handle small messages coming from the network.
@@
As a second example: for throughput-oriented workloads DMA has evolved to efficiently transfer data to and from main memory without polluting the CPU cache. However, for small, fine-grained interactions, it is important that almost all the data gets into the right CPU cache as quickly as possible.
@@
These reasons led us to rethink the DMA-based model, and whether directly involving the CPU in data transfers to and from devices might be a better approach for these scenarios on modern hardware. We are not the first to suggest this direction for some cases (Larsen and Ben Lee, 2015; Chaudhari et al., 2017), but we push the question much further: given a future cache-coherent interconnect, what new ways of using Programmed I/O might become compelling for data center applications?
Note that we are not proposing abandoning the traditional model, or some of its enhancements we survey in Section 7, for throughput-oriented applications, which can continue to ride the current hardware curve for bandwidth. Instead, we argue it is time to examine complementary alternatives.
@@ 3. Experimental platform
Figure 1. XDMA invocation latency comparison.
Figure 2. PIO invocation latency comparison.
We investigate the modern trade-offs between PIO, DMA, and cache-coherent interconnects using a real hardware platform rather than a simulator. Simulation would be appropriate for detailed quantitative comparisons of known techniques, but in this paper we explore unconventional uses of cache coherence and so want to establish not only that our techniques are performant, but that they can practically be implemented in a real system.
The Enzian research computer (Cock et al., 2022) can be thought of as a two-socket NUMA server, where one socket houses a Marvel Cavium ThunderX-1 CPU, with a Xilinx XCVU9P Field Programmable Gate Array (FPGA) in the other. The CPU is a 48-core ARMv8 processor running at 2.0GHz, and 128GiB of 2133MT/s DDR4 memory spread across 4 memory controllers. Each core has a 32KiB, 32-way associative, write-through, physically indexed, physically tagged, L1 data cache. These are connected to a 16MiB, write-back, 16-way associative, shared L2 last-level cache. Hardware keeps the L1 caches coherent with the L2, and so in ARM architecture parlance the point of coherence is the L1 cache, whereas the point of unification is the L2 cache write-through. The cache-line size is 128 bytes – note that this is double the conventional 64-byte line size.
The CPU and FPGA are connected by the CPU’s native inter-socket cache coherent protocol, the Enzian Coherence Interface (ECI), which Enzian implements on the FPGA. It is a MOESI-like directory protocol with the physical address space statically partitioned between the NUMA nodes (CPU and FPGA). It uses 24 bidirectional lanes (organized into 2 12-lane links), each running at about 10Gb/s, for a theoretical inter-socket bandwidth of about 30GiB/s.
@@
While Enzian provides a real hardware platform with server-class performance, direct comparison with modern server platforms is not straightforward: the latter do not offer full cache coherence between CPU and devices, but have more recent, faster cores, memory systems, and PCIe.
@@
We therefore calibrate against a modern PC for both PIO and DMA over PCIe with benchmark comparisons of Enzian and a modern PC (Intel Core i7-7700 3.6GHz Kaby Lake, PCIe Gen.3 x16) with 64-byte cache lines connected to an AMD Virtex UltraScale+ VCU118 card, using the “write then read” experiment detailed in Section 6.1. The VCU118 uses the same FPGA as Enzian, albeit a slightly slower speed gauge.
DMA performance over PCIe:
@@
The first experiment uses the Xilinx XDMA IP and its descriptor-based protocol to transfer data between CPU and FPGA over PCIe. Figure 1 shows results for various data transfer sizes, running the Xilinx driver in both interrupt-driven and polled modes.
@@
We see that XDMA over PCIe is about 3 times faster on the PC than on Enzian, while the difference between interrupt-driven and polling performance is much insignificant. Single transaction latency is almost constant on both platforms up to the PCIe transaction size limit of 4KiB, and then increases.
@@
The performance difference between the machines is due to several factors. For small transfers, the CPU overhead of descriptor setup dominates, and an x86 core is simply faster than a ThunderX-1 core. In addition, the memory system is somewhat faster on the PC. The factor of 2 difference in PCIe bandwidth between the two platforms does not appear to be a factor in these experiments.
PIO performance over PCIe
@@
In the second calibration experiment the CPU performs PIO reads and writes over PCIe to the FPGA. We experimented with several different optimizations in this experiment (see Section 4), in particular the use of write combining and vector instructions. We present the best figures obtainable on each platform.
@@
Figure 2 shows that for transaction sizes above 32 bytes, the PC is about twice as fast as Enzian over the PCIe cable between CPU and FPGA. Here, the difference in PCIe bandwidth is responsible for the latency difference, in particular for reads (limited to 128 bits on both platforms). Writes, in contrast, are combined and pipelined at the PCIe interface.
4. Programmed I/O over PCIe
PCIe is complex, and we spent considerable time optimizing CPU-initiated reads and writes for different platforms. We describe here how we arrived at optimally-efficient PIO latency to the FPGA over PCIe.
We pre-map PCIe apertures (BARs) into user space on the CPU and measure latency as the time to write a value to this space (thereby invoking a function on the device) and read a result back (Figure 3).
// map 4KiB of PCIe memory for read/write
int fd = open("/sys/bus/pci/devices/.../resource0",
    O_RDWR);
void *pcie_bar = mmap(NULL, 4096,
    PROT_READ | PROT_WRITE, MAP_SHARED, fd, 0);
// write arguments to PCIe memory
*(volatile invoke_args_t *)pcie_bar = invoke_args;
__sync_synchronize();
// read result from PCIe memory
invoke_res_t invoke_res = *(volatile invoke_res_t *)pcie_bar;
__sync_synchronize();
Figure 3. An invocation using PIO & PCIe. Error handling omitted.
PCIe memory writes are posted transactions and may complete out-of-order, allowing multiple requests in-flight at the same time. Most systems also have dedicated bus units to combine writes and coalesce transactions. Enzian’s ThunderX-1 performs write-combining for stores to PCIe, allowing 512 bits per bus round-trip (confirmed with a logic analyzer) and resulting in efficient PIO write transactions.
[…]
This means that on ThunderX-1, common optimization techniques like vector instructions have negligible benefit for PIO over PCIe: the hardware already coalesces writes into 512-bit transactions and reads are limited by the read bus width of 128 bits.
x86 has more optimization opportunities for PIO reads and writes. The Intel Data Streaming Accelerator (Intel, 2024) offers MOVDIR64B and ENQCMD instructions, allowing 64-byte PIO writes with write-combining mappings and user-space descriptor submission without driver intervention. Mapping memory with write-through attributes, as suggested by Tide (Humphries et al., 2024), allows cache line-sized PIO reads. These architecture-specific extensions are not available on ThunderX-1.
5. PIO over a coherent interconnect
@@
We now turn to programmed I/O over a cache coherent interconnect. A naive approach would have the CPU simply copy data over a coherent interconnect as in conventional PIO. This requires careful cache management on the CPU using explicit cache flush and invalidate instructions, but is appropriate when the hardware standard mandates a cache on the device, as with OpenCAPI 4.0 (Stuecheli et al., 2018; Allison et al., 2020).
@@
However, we are more interested in how a coherent interconnect can best perform low-latency, efficient data movement between CPU registers and device hardware, including cases where the conventional notion of the cache as a view over memory contents is not appropriate, such as data arriving over a network.
Scoping:
@@
We scope the scenario further to the case where the CPU cores, or their Last-Level Cache (LLC), communicate with devices via a MESI-like (Nagarajan et al., 2020), message-based coherence protocol with stable states corresponding to Modified (dirty), Shared (clean), Exclusive (clean), and Invalid, using a distributed cache directory. Note that this means that each address or cache line has a home node, which maintains the directory entry for the line.
@@
We also assume a symmetric protocol: each device is a first-class participant in the protocol with a subset of physical addresses homed at it, and the device therefore maintains directory state for these lines (although it might not actually cache data itself). Moreover, we take the perspective of the device designer: the device sees individual cache messages, rather than transparently coherent memory.
@@
These assumptions are rare in hardware today but they do hold for proposed standards like CXL 3.0 (CXL Consortium, 2024), and also for the Enzian platform we use in experiments.
Baseline: simple coherence
@@
We first discuss using simple cache coherence to re-apply existing techniques inter-core communication in software, such as FastForward (Giacomoni et al., 2008). Here (Figure 4), the sender (CPU) writes data to a cache line on which the receiver is spinning reading. Assume the line starts in Shared state in both caches. When the sender writes the line (typically atomically via the core’s write buffer), it transitions briefly to Modified state, before being fetched by the receiver over the interconnect and reverting to Shared.
@@
This can be highly efficient in software, particularly when deployed as a ring buffer of cache lines, and results in minimal interconnect traffic (since the receiver spins on a local copy of the line), although the CPU itself must spin. Indeed, this is the technique recently used in CC-NIC (Schuh et al., 2024) to emulate a NIC using a different CPU socket.
However, this design assumes that each endpoint can only manipulate its local cache using software instructions (loads, stores, etc.) and also that flush or invalidate instructions should be avoided for performance reasons. This is not the case when one of the endpoints is a device with message-level access to the cache coherence protocol, and possibly maintaining its own directory state.
The implications of message-level access
A device on a symmetric coherent interconnect can see more cache state and events, and also initiate more operations, than a CPU core can. Note that this does not entail a cache per se on the device at all, indeed this is rarely appropriate.
First, the device receives messages from the CPU cache that request lines in Shared or Exclusive state, or which request that the caching state on the device be downgraded (e.g., from Exclusive to Shared, or from Shared to Invalid). It is free to react internally to these events as it chooses, as long as it does not violate the protocol.
@@
Fourth, unlike a conventional cache, the device does not have to respond immediately to every request from the CPU’s cache. Instead, it can choose to delay the response (blocking the requesting core) until some other operations complete. Care must be taken to respond before any hardware-imposed timeouts, but these are typically generous (hundreds of milliseconds on the ThunderX-1, for example) and can be worked around in software (e.g., by sending a “try again” response within the timeout).
@@
Communication protocol design variants
We present 4 different but similar communication protocols (shown in Figure 5) which vary based on whether (a) the CPU or the device initiates the message transfer, and (b) the cache lines used are homed on the CPU or device. All provide a synchronous invocation: a transfer of a single cache-line-sized message, followed by a single cache-line response. All use 2 cache lines, which we designate A and B. These protocols can be generalized to larger payloads, and latency will grow linearly with the payload size. Interestingly, they can be made slightly more efficient by using more cache lines, but quickly become complex. We focus on the simple 2-line case in this paper.
The protocols use the same techniques and insights, and are best described using the example of a CPU-initiated invocation on the accelerator using addresses homed on the device (5(a)). An implementation of this protocol is used for the evaluation in Section 6.1. Note that, unlike RPC, invocations are local to the machine and avoid marshaling or unmarshaling of messages.
@@
    The protocol starts with line B in Exclusive state in the CPU cache, and line A Invalid (we show how this can be conveniently achieved below). To initiate an operation, the CPU writes the request data from its registers into line B, which transitions to Modified. This happens without communication with the device, since B was previously in Exclusive.
@@
    The CPU cache receives line A’s new data and the CPU unblocks, immediately loading the first word of the invocation response into a register.
As described, the end state is that A is now Invalid and B is Shared in the CPU cache. However, in some deployed MESI-like protocols (including the Enzian interconnect we use), the device is free to return the data stipulating that it is not Shared but now Exclusive to the CPU. This final feature leaves the cache in the state it started in, with the roles of A and B reversed, ready for the next transaction.
The whole exchange takes 2 round trips on the coherent interconnect, and require no spinning on the CPU.
The other 3 protocols work similarly. For example, 5(b) shows the analogous case where the device initiates the request to the CPU, but still using lines homed on the device. This is used for the NIC experiment in Section 6.3.
@@
Subtleties:
A number of issues must be handled to make this practical, starting with timeouts: we block a core’s request for a cache line until a device operation has completed. A timeout in the CPU’s cache is likely to cause a memory fault exception, a fatal machine check, or the processor simply locking up (on Enzian, it is a machine check).
We solve this issue with a small state machine on the device that returns a “not ready yet” response before an impending timeout, causing the software on the issue another request for the other cache line, extending the response time indefinitely without the need for spinning. In practice, we target operations that are shorter than the timeout.
@@
The second issue concerns deadlock due to the device stalling a request until it can issue, and get a reply back for, another request to a different cache line. Existing standards for interoperable cache coherent interconnects are silent on what happens in this situation, possibly because it has not occurred until now. The design assumes that transactions on different cache lines can progress independently.